# Toward a Physical Theory of Intelligence

### v1

@@ Abstract
We present a physical theory of intelligence grounded in irreversible information processing in systems constrained by conservation laws. An intelligent system is modelled as a coupled agent–environment process whose evolution transforms information into goal-directed work. To connect information to physical state, we introduce the Conservation–Congruent Encoding (CCE) framework, in which encodings correspond to metastable basins of attraction whose separability is enforced by conservation laws. Within this framework, intelligence is defined as the amount of goal-directed work produced per nat of irreversibly processed information. From this definition we derive a hierarchy of physical constraints governing information intake, irreversible computation, and work extraction in open systems. The framework reveals how long-horizon efficiency requires the preservation of internal informational structure, giving rise to self-modelling, and it establishes that physically embodied intelligent systems possess intrinsic epistemic limits analogous to incompleteness phenomena. Applying the theory to biological systems, we analyse how oscillatory and near-critical dynamics optimise the trade-off between information preservation, dissipation, and useful work, placing the brain near an efficient operating regime predicted by the framework. At the architectural level, we develop a theory of continuous dynamical circuits in which classical Boolean logic emerges as a special case of attractor selection, while more general invariant geometries support computational modes beyond fixed-point logic. Finally, we propose a physically grounded perspective on artificial intelligence safety based on irreversible information flow and structural homeostasis. Together, these results provide a unified, substrate-neutral account of intelligence as a physical phenomenon.
@@ 1 Introduction
Intelligence is typically studied through behavioural, information-theoretic or algorithmic lenses, yet all intelligent systems are ultimately physical systems. They evolve in time, consume resources, store and transform information, and interact with their environments through structured flows of energy and matter. Despite this, there is no unified account linking physical laws to the computational properties associated with intelligence. Existing accounts rely primarily on behavioural benchmarks or information-theoretic abstractions and therefore lack an explicit grounding in physical law, making principled comparison across heterogeneous systems and the derivation of fundamental limits difficult.
This paper develops a physical theory of intelligence grounded in the dynamics of conserved quantities and information. We model an agent and its environment as coupled dynamical systems exchanging information through well-defined ports. To connect encoded information to physical configuration, we introduce the Conservation-Congruent Encoding (CCE) framework, under which encodings correspond to metastable basins of attraction whose separation is enforced by conservation laws. CCE provides a substrate-neutral bridge between information and physical descriptions, allowing information processing to be tied directly to physical units.
Within this framework, we define the intelligence of a physical system as the rate at which it converts irreversibly processed information into goal-directed work. This definition is grounded in the reversible–dissipative (metriplectic) decomposition of nonequilibrium dynamics, which separates reversible, structure-preserving flows from dissipative, entropy-producing transformations.
The theory yields two further consequences. First, self-modelling and consciousness emerge as the persistent informational structures an agent must maintain in order to reduce irreversible computation over long horizons. We formalise this intuition by defining a measure of consciousness dual to intelligence and prove that any physically embodied agent with finite preserved-information capacity is necessarily epistemically incomplete; it cannot form a fully faithful model of its own internal dynamics. Second, the framework predicts that certain classes of dynamics, most notably oscillatory and near-critical regimes, optimise the tradeoff between information preservation, irreversible information processing, and goal-directed work. We show that these regimes, widely observed in biological neural systems, correspond to locally optimal operating points under the proposed theory.
Beyond analysis, we construct a compositional theory of continuous dynamical circuits using ports and attractor semantics under CCE. We show that classical Boolean logic arises as a special case of fixed-point computations and demonstrate how more general attractor geometries support computational modes unavailable to digital logic. A detailed case study of oscillator-based frequency discrimination illustrates how geometry-aligned dynamics achieve the same behavioural objective with orders-of-magnitude less irreversible information processing than a digital implementation.
Finally, we derive physically motivated constraints on irreversible information flow, entropy production, and structural homeostasis, yielding intrinsic safety guardrails for the development and deployment of engineered intelligent systems. These constraints embed safety directly into the physical dynamics of an agent, ensuring operation remains within physically sustainable regions of state space. We further outline future research directions for artificial intelligence safety centered on the coupled dynamics of intelligent systems.
Taken together, these results provide a unified, substrate-neutral account of intelligence as a physical phenomenon. By linking reversible and irreversible information processing to goal-directed work, the theory offers a common language for analysing biological, artificial, and composite intelligent systems and establishes a foundation for designing future systems within the limits imposed by physical law.
@@ 2 Mathematical Framework for Assessing Intelligence
This section formalizes the physical coupling between an agent and its environment, providing the foundation on which intelligence is defined as a measurable property of their interaction. Both the agent and the environment are modelled as stochastic dynamical systems whose states evolve over time and are coupled through explicit observation and action ports. These ports constitute the physical interface through which goal-directed behaviour emerges.
The framework is developed in four stages:
@@
    representation of goals as energy-based functionals over trajectories;
@@
    introduction of a physically grounded information-encoding scheme that connects computational irreversibility to energetic cost;
@@
    definition of intelligence as a scalar measure derived from these structural elements.
2.1 Agent-Environment Dynamics Model
Let $(X,\mathcal{X})$ and $(E,\mathcal{E})$ denote the measurable state spaces of the internal (agent) and external (environment) systems, respectively. Define measurable port maps
$\Gamma_{X}:X\to\mathscr{X},\qquad\Gamma_{E}:E\to\mathscr{E},$ (1)
where $\mathscr{X}\subseteq X$ and $\mathscr{E}\subseteq E$ are the agent and environment interface subspaces. The coupled one-step transition dynamics are defined by measurable kernels
$X_{t+1}$ $\sim K_{X}(\,\cdot\,\mid X_{t},\,\Gamma_{E}(E_{t})),$ (2)
$E_{t+1}$ $\sim K_{E}(\,\cdot\,\mid E_{t},\,\Gamma_{X}(X_{t})),$ (3)
@@ 2.1 Agent-Environment Dynamics Model
$X_{t+1}$ $\perp\!\!\!\perp E_{t}\mid(X_{t},\,\Gamma_{E}(E_{t})),$ (4)
$E_{t+1}$ $\perp\!\!\!\perp X_{t}\mid(E_{t},\,\Gamma_{X}(X_{t})).$ (5)
Hence, the ports $(\Gamma_{X}(X_{t}),\,\Gamma_{E}(E_{t}))$ form a Markov blanket that separates internal and external states. Let $S_{t}:=(X_{t},E_{t})$ denote the joint process with state space $(X\times E,\;\mathcal{X}\otimes\mathcal{E})$, and let $\mu_{t}:=\mathsf{Law}(S_{t})\in\mathcal{P}(X\times E)$ denote its probability law. The joint transition kernel $K$ on $(\mathcal{X}\otimes\mathcal{E})$ factorizes via the ports as
$K(dx^{\prime},de^{\prime}\mid x,e)$ $=K_{X}(dx^{\prime}\mid x,\,\Gamma_{E}(e))\;K_{E}(de^{\prime}\mid e,\,\Gamma_{X}(x)).$ (6)
@@
which defines a valid measure-valued dynamical system on $(X\times E,\;\mathcal{X}\otimes\mathcal{E})$. Under standard regularity assumptions (standard Borel spaces, measurable kernels, and optional Feller continuity), this recursion guarantees the existence of a unique probability law for the process $\{S_{t}\}_{t\geq 0}$ given an initial distribution $\mu_{0}$. The pushforward
$\bar{\mu}_{t}:=(\Gamma_{X},\Gamma_{E})_{\#}\mu_{t}$ (8)

### v2

@@ Abstract
While often treated as abstract algorithmic properties, intelligence and computation are ultimately physical processes constrained by conservation laws. We introduce the Conservation-Congruent Encoding (CCE) framework as a unified, substrate-neutral physical framework for studying intelligence. We propose that information processing emerges when open systems undergo irreversible transitions, carving out macroscopic states from underlying reversible micro-dynamics. Generalizing Landauer’s principle to arbitrary conserved quantities via metriplectic flows, we derive a universal bound for macroscopic computation. This yields physical metrics for intelligence and an operational analogue for consciousness, quantifying an agent’s ability to extract work from the environment while minimizing its own dissipative dynamics. Applying CCE to the limits of physical observation, we model measurement as an active coarse-graining process rather than a passive projection. At the quantum scale, CCE recovers the Lindblad Master Equation, consistent with modelling decoherence as the dissipative exhaust required to record a measurement. Scaling to cosmological limits, we explore the hypothesis that gravity emerges as the macroscopic geometric footprint of these bounds. We show that, under this hypothesis, measurement-induced dissipation is consistent with a volumetric phase-space collapse, offering a dynamical route to the Bekenstein-Hawking area law. Equating the Landauer exhaust of this coarse-graining to horizon deformation outlines a limiting-case recovery of the Einstein Field Equations. Ultimately, by establishing a substrate-neutral link between thermodynamic dissipation, quantum measurement, and spacetime geometry, CCE provides physical constraints for understanding both natural and artificial intelligence.
@@ 1 Introduction
Intelligence is typically studied through behavioural, information-theoretic, or algorithmic lenses, yet all intelligent systems are ultimately physical systems. They evolve in time, consume resources, store and transform information, and interact with their environments through structured flows of energy and matter. Despite this, there is no unified account linking physical laws to the computational properties associated with intelligence. Existing accounts rely primarily on behavioural benchmarks or information-theoretic abstractions, and therefore lack an explicit grounding in physical law, making principled comparison across heterogeneous systems and the derivation of fundamental limits difficult. Furthermore, existing physical theories have historically remained largely confined to thermodynamics, linking computation to the export of heat rather than unifying concepts across all conserved quantities. Consequently, fields such as neuroscience, artificial intelligence, and quantum mechanics have developed disparate theoretical languages to describe how systems encode memory, process information, and exert causal influence over their environments.
This paper proposes a unified physical framework for studying intelligence, grounded in the dynamics of conserved quantities and irreversible information processing. We connect abstract computation to explicit physical costs by introducing the Conservation-Congruent Encoding (CCE) condition. Under CCE, informational distinctions correspond to metastable basins of attraction whose separability is enforced by fundamental conservation laws. Crucially, we show that the irreversible formation or collapse of these dynamically distinguishable equivalence classes requires the export of a conserved quantity, thereby generalising Landauer’s bound across arbitrary physical domains. Whether an agent encodes information in the chemical potential of biological wetware, the electrical charge of a silicon GPU, or the angular momentum of a quantum spin, we show that the minimal energetic cost of irreversible computation is governed by the product of the channel’s conjugate force ($\mathcal{F}$) and its fundamental characteristic scale ($\alpha$).
Within this physically grounded framework, we model an agent and its environment as coupled dynamical systems. We represent goals by macroscopic work performed over spatiotemporal trajectories. We define two complementary, scale- and boundary-relative physical efficiencies: Intelligence ($\chi$) measures the macroscopic goal-directed work produced per nat of irreversibly processed information, while consciousness ($\kappa$)—framed here strictly as an operational property—measures the goal-directed work supported per nat of preserved internal information. Furthermore, rather than assuming fixed system architectures, we treat the specification of the system boundary itself as a dynamic optimization problem, identifying the physically meaningful agent as the minimal boundary that maximises these efficiencies.
To rigorously validate the substrate neutrality of the CCE framework, we stress-test it at both the microscopic and macroscopic limits of physical observation. At the quantum scale, by framing the double-slit experiment as an information-extraction problem requiring the physical instantiation of a CCE, we map our generalised metriplectic dynamic equations directly to Hilbert space. This recovers the Lindblad Master Equation, suggesting that the destruction of quantum interference can be modelled as a consequence of the generalised dissipative exhaust required by a macroscopic detector to record a classical bit.
Scaling to the cosmological limit, we then explore the hypothesis that gravity emerges as the macroscopic geometric footprint of these informational bounds. We address the fundamental geometric paradox between volume-preserving microscopic dynamics and area-bounded macroscopic limits by showing that, under this hypothesis, measurement-induced dissipation is consistent with a volumetric phase-space collapse. By equating the localised Landauer exhaust of this coarse-graining to causal horizon deformation, we recover the Einstein Field Equations in this limiting construction via the Raychaudhuri integration. Extending this to the extreme limits of physical measurement, we establish a strict polynomial bound on macroscopic observation, suggesting that the structural fuel required for an autonomous observer to record a black hole’s microstate may trigger mutual gravitational collapse if the observer crosses a critical threshold of approximately $2.52R_{S}$.
Beyond these core definitions, we extend these metrics to coupled multi-component systems, yielding a quantitative notion of emergent intelligence as a coupling-dependent change in efficiency across scales. We analyse how reversible structure-preserving dynamics can reduce the irreversible information required for sustained goal-directed work, with oscillatory and near-critical regimes providing illustrative biological examples. We also develop a compositional theory of continuous dynamical circuits using ports and attractor semantics under CCE, and outline a physics-grounded perspective on AI safety based on strict constraints on irreversible information flow, entropy production, and structural homeostasis.
Taken together, the results propose a substrate-neutral theory that seeks to relate notions of intelligence directly to fundamental physical law, making its constraints and trade-offs explicit across all scales of nature.
@@ 2 Mathematical Framework for Assessing Intelligence
This section develops a mathematical framework that renders intelligence a physically measurable property of coupled agent–environment dynamics. The central objective is to distinguish information transport from information use, and to identify when informational transformations incur unavoidable physical cost. To this end, both the agent and the environment are modelled as stochastic dynamical systems whose states evolve in time and are coupled through explicit ports. These ports define the physical interface across which information is exchanged and goal-directed behaviour emerges.
The framework isolates the minimal structural elements required to quantify intelligence as a physically grounded efficiency. It is developed in four stages:
@@
    definition of useful work completed by the agent under a spatiotemporal framing;
@@
    introduction of a physically grounded information formulation that links computational irreversibility to physical cost;
@@
    definition of scale- and boundary-relative intelligence metrics derived from these elements.
2.1 Agent–Environment Dynamics Model
Let $(X,\mathscr{X})$ and $(E,\mathscr{E})$ denote the measurable state spaces of the internal (agent) and external (environment) systems, respectively. Define measurable port maps
$\Pi_{X}:X\to\Xi_{X},\qquad\Pi_{E}:E\to\Xi_{E},$ (1)
where $\Xi_{X}$ and $\Xi_{E}$ are measurable port spaces representing the agent–environment interface variables at the chosen system boundary and scale. The coupled one-step transition dynamics are defined by measurable kernels
$X_{t+1}$ $\sim K_{X}(\,\cdot\,\mid X_{t},\,\Pi_{E}(E_{t})),$ (2)
$E_{t+1}$ $\sim K_{E}(\,\cdot\,\mid E_{t},\,\Pi_{X}(X_{t})),$ (3)
@@ 2.1 Agent–Environment Dynamics Model
$X_{t+1}$ $\perp\!\!\!\perp E_{t}\mid(X_{t},\,\Pi_{E}(E_{t})),$ (4)
$E_{t+1}$ $\perp\!\!\!\perp X_{t}\mid(E_{t},\,\Pi_{X}(X_{t})).$ (5)
Hence, the ports $(\Pi_{X}(X_{t}),\,\Pi_{E}(E_{t}))$ form a Markov blanket that separates internal and external states, following the formalisms established in graphical models [18] and subsequently applied to the variational foundations of self-organising systems [10, 16]. Let $S_{t}:=(X_{t},E_{t})$ denote the joint process with state space $(X\times E,\;\mathscr{X}\otimes\mathscr{E})$, and let $\mu_{t}:=\mathsf{Law}(S_{t})\in\mathcal{P}(X\times E)$ denote its probability law. The joint transition kernel $K$ on $(\mathscr{X}\otimes\mathscr{E})$ factorises via the ports as
$K(dx^{\prime},de^{\prime}\mid x,e)$ $=K_{X}(dx^{\prime}\mid x,\,\Pi_{E}(e))\;K_{E}(de^{\prime}\mid e,\,\Pi_{X}(x)).$ (6)
@@
which defines a valid measure-valued dynamical system on $(X\times E,\;\mathscr{X}\otimes\mathscr{E})$. Under standard regularity assumptions (standard Borel spaces, measurable kernels, and optional Feller continuity), this recursion guarantees the existence of a unique probability law for the process $\{S_{t}\}_{t\geq 0}$ given an initial distribution $\mu_{0}$. The pushforward
$\bar{\mu}_{t}:=(\Pi_{X},\Pi_{E})_{\#}\mu_{t}$ (8)